How do I insert parentheses to make these statements true? Thank you!
-8+3-2+7=-2
-8+3-2+7=-16

Answer by josmiceli(6797)

My first observation is, if you had all plusses like:
1 + 2 + 5 + 6 + 4 + 1 + 9
then it wouldn't matter where you put the parentheses.It wouldn't change anything.
The same is true for all minuses, like:
- 3 - 5 - 6 - 1
so, you can only change things where minus before plus, or a plus before minus.
-(8 + 3 - 2) + 7 = -(9) + 7 = -2
next one:
does -(8 + 3) - (2 + 7) = -16 ?
no, it equals -20
I finally came up with:
-(8 + 3 - 2 + 7) =
-(11 - 2 + 7) =
-(11 + 5) = -16
and that works.